# These $\altaffiliationThese authors contributed equally.$

### v1

@@ 2.2.1 Preprocess the Input File
Figure 1: Preprocessing of the input test.zip. (a) User-provided input file and message. (b) Generation of the preprocessing Python script using tool calling and the provided Python function templates. (c) Final preprocess.py and the resulting XYZ files. (d) Analysis of molecular structures with the visual Mentor, as well as a image of the molecular structure and atomic labels. (d) Analysis of molecular structures by the Mentor, including the molecular structure and atomic labels of keto isocytosine shown in mol.png. (e) Six generated descriptor sets based on RICs.
@@
Clearly, as with output files generated by other NAMD packages, the raw JADE data require extensive preprocessing before formal analysis. This preprocessing includes unzipping all output files, validating their contents, collecting the initial geometries, identifying the hopping times, and extracting the corresponding hopping geometries. Here, we emphasize that although the workflow and performance of VisU are demonstrated using keto isocytosine based on JADE outputs, the overall data-processing pipeline of VisU is, in principle, directly applicable to outputs from other NAMD codes without further modification.
@@
Rather than directly generating a complete preprocessing script in a single step, the Engineer first reasoned about the information required for subsequent analysis, including the directory structure, file naming conventions, and the locations of trajectory and geometry files. Guided by this reasoning, it selectively invoked file-system inspection tools embedded within VisU to explore the unzipped archive and extract the necessary structural information. After acquiring the workspace context, the Engineer composed a preprocessing script from a curated library of expert-validated function templates. In this sense, the overall preprocessing procedure closely resembles human expert operations. The Student then executed the script and returned error messages in cases of failure. In the present example, the program encountered a Python error (FileNotFoundError) during the first execution. The Student collected all error information and relayed it to the Engineer, which subsequently revised the script. The data-preprocessing task was completed successfully after a second iteration. Details of the error analysis and the corrected preprocessing script are provided in the SI. As shown in the previous section, this preprocessing step produces three required output files, namely ref.xyz, init.xyz, and hopping.xyz (Figure 1.(c)).
@@
Based on the reference geometry (ref.xyz) and all hopping geomeries (hopping.xyz), VisU generates the molecular descriptors. The Student plotted a 2D molecular image (mol.png) with atomic labels in the basis of the reference geometry, as illustrated in Figure 1.(d). This image is then passed to the Mentor, powered by the Doubao-Seed-1.6-Vision model, which was carefully prompted to understand the context of the analysis workflow.
Leveraging its visual-reasoning capability, this Mentor first “sees” the image and determines whether the input molecule is compatible with the current available protocol. For the present system, VisU returns the list of atoms constituting the ring moiety, which serves as explicit guidance for the Student to construct the corresponding RIC descriptors. Using the hopping geometries collected in hopping.xyz, the Student then built six sets of RIC descriptors, as shown in Figure 1.(e): $D_{\mathrm{ring}}$, $A_{\mathrm{ring}}$, $R_{\mathrm{ring}}$, $D_{\mathrm{eg}}$, $A_{\mathrm{eg}}$, and $R_{\mathrm{eg}}$. These descriptors provide the foundation for subsequent unsupervised channel discovery and mechanistic analysis. In parallel, the same descriptor sets are computed for all initial geometries in init.xyz.
@@ 2.2.2 Extract the Channels
Figure 2: “Mentor-Student” cooperation during the first iteration of unsupervised channel discovery. Based on diagnostic plots produced by the Student, the carefully prompted visual-language Mentor (Doubao-Seed-1.6-Vision) provides targeted guidance on: (a) selection of the optimal reduced-dimensional subspace, (b) choice of clustering algorithm, (c) hyperparameter tuning, and (d) final evaluation of the resulting clusters.
@@
Using the prepared descriptor sets, channel discovery proceeds via an iterative PCA-then-clustering analysis protocol zhuPrincipalComponentAnalysis2022a.
@@
As shown in Figure 2.(a), the first iteration of channel discovery begins by applying PCA to each of the six descriptor sets to obtain low-dimensional representations of the hopping geometries. In all cases, the first two principal components explain more than 60% of the total variance. The resulting PCA projections are then forwarded to the Mentor for assessment. In the present example, the Mentor leverages its visual-reasoning capability to identify that the hopping points are well separated in the $D_{\mathrm{ring}}$-based reduced space, indicating the presence of several distinct channels and highlighting the important role of ring-associated dihedral motions in the nonadiabatic dynamics.
@@
The next critical step is to group similar hopping geometries into clusters. In Figure 2.(b), the Mentor inspects the distribution in the $D_{\mathrm{ring}}$ PCA plot using its visual-reasoning capability and recommends agglomerative clustering for the initial partitioning. As shown in Figure 2.(c),the Student first performs clustering with the default parameter $k=2$. Upon reviewing the resulting plot, the Mentor immediately identifies that the central region is visually distinct and should not be merged with the adjacent cluster. Therefore, the Mentor instructs the Student to adjust the parameter to $k=3$, after which the clustering result aligns with expectations. Following re-clustering with the corrected parameter, the Mentor confirms that the current clustering is accurate.
@@
A final visual inspection by Mentor, shown in Figure 2.(d), verifies that the current clustering result with three clusters is reasonable, ensuring that this step aligns with expert-level visual intuition. The first discovery iteration thus concludes with three clusters: Cluster 0, Cluster 1, and Cluster 2. Among them, Cluster 2 contains only 37 structures, corresponding to less than 15% of the total dataset. This cluster is therefore treated as an independent candidate channel (Candidate Channel 0), while the remaining two clusters proceed to a second iteration of analysis. Afterward, the same PCA-then-clustering procedure is then recursively applied to all remaining subsets. As illustrated in Figure S1, all clusters either become non-separable after three iterations or contain fewer than 15% of the structures, ultimately yielding seven nonadiabatic decay channels (Candidate Channel 0 to 6).
@@
As different data patterns appear in these iterative steps, distinct clustering algorithms are recommended based on their distributions: including agglomerative clustering (Iteration 1), DBSCAN (Iteration 2), and K-Means (Iteration 3), as in Figure S1. Additional results are provided in the SI. These observations demonstrate that VisU can effectively handle both algorithms like agglomerative clustering, which requires only a single hyperparameter and aligns closely with visual intuition, and more complex methods such as DBSCAN, where the two parameters (eps and min_samples) are less directly interpretable from visual inspection.
@@ 2.2.3 Identify the Important Molecular Motion
At this stage, the Student performed PCA projections of both the initial and hopping geometries, submitting the resulting two-dimensional representations to the Mentor for analysis. By visually inspecting these projection plots, the Mentor determines which sets of RICs are most effective for distinguishing hopping from initial geometries. In the present example (Figure 3.(a)), the hopping and initial geometries are clearly separated in the reduced space defined by $D_{\mathrm{ring}}$, $D_{\mathrm{eg}}$, $A_{\mathrm{ring}}$, and $R_{\mathrm{ring}}$, indicating that these RIC sets capture the leading reactive coordinates in the photochemical processes.
@@ 2.2.4 Validate and Summarize Mechanisms
Figure 4: (a) Final nonadiabatic channel analysis tree. (b) Averaged and (c) representative structures (hydrogen atoms omitted) for the four channels.
@@
An essential task for the Mentor is to check for chirality. For example, several candidate channels in the present system were identified as mirror counterparts based on both qualitative visual inspection and quantitative data comparison. As shown in Figure 4.(a), the Mentor merged different groups showing mirror-symmetry: Candidate Channel 1 and Candidate Channel 6 into Channel B; Candidate Channel 3 and Candidate Channel 5 into Channel C; and Candidate Channel 2 and Candidate Channel 4 into Channel D. Candidate Channel 0 exhibited no mirror-symmetric counterpart and was therefore retained as Channel A. As in Figures 4.(b) and (c), summarize the initial seven candidate channels were finally consolidated into four distinct nonadiabatic channels: Channel A, B, C, and D.
It is important to note that different decay channels may be governed by the same nuclear motions but with varying magnitudes. For instance, Channel B, C, and D all exhibit pronounced out-of-plane motions of O9 and C10, while their direction differ. In principle, one could perform systematic pairwise comparisons of initial and hopping geometries across channels; however, this approach becomes tedious in multichannel situations. Instead, VisU relies on the Mentor to perform a holistic comparison by integrating key features, representative geometries, and dominant motions across all channels. Through direct visual inspection and cross-channel assessment, the Mentor clarifies the mechanistic similarities and differences among the nonadiabatic decay channels.
Finally, the Mentor draws the following conclusions. Channel A accounted for approximately 9% of the total hopping events, which was characterized by a dominant C9-C10 stretching motion. In contrast, the remaining three channels involved combinations of C1 puckering and O9-C10 out-of-plane motions. Among them, Channel B, Channel C, and Channel D contributed approximately 10%, 6%, and 75% of the total population, respectively. Although these three channels share similar core motions, they differ in the directionality of the side-group motion associated with the O9-C10 fragment, as shown in Figures 4.(b) and (c). These results are consistent with our previous and recent studies zhuPrincipalComponentAnalysis2022a, liu2025automatedframeworkanalyzingstructural, with additional discussion provided in the SI.
At the end, the complete mechanistic interpretation, including channel populations, representative structures, and visual summaries derived from the raw nonadiabatic dynamics trajectories, is compiled in a comprehensive analysis report provided in the SI. With this, the end-to-end analysis performed by VisU is complete, where was executed fully automatically without human intervention.
@@ 3 Conclusion
The powerful analysis capability of VisU is stems from the integration of two state-of-the-art large language models: Doubao-Seed-1.6-Vision, which provides expert-level visual reasoning, and DeepSeek-V3.2, which enables adaptive tool-calling and script construction. This synergy facilitates a multi-agent paradigm that orchestrates a comprehensive four-module workflow: Preprocessing, Recursive Channel Discovery, Important-Motion Identification, and Validation and Mechanistic Summary.
In a representative case study, VisU successfully identified four distinct nonadiabatic channels, uncovered the corresponding leading reactive coordinates, and automatically handled mirror-symmetry merging. Furthermore, it performed cross-channel comparisons and generated a comprehensive, figure-rich academic report without human intervention. The framework demonstrated a professional-grade ability to reproduce dominant mechanistic features reported in previous work zhuPrincipalComponentAnalysis2022a, liu2025automatedframeworkanalyzingstructural. Remarkably, starting only from raw NAMD simulation outputs and minimal text prompts, the entire protocol is executed through autonomous visual inspection and expert-like guidance, powered by advanced vision-reasoning LLMs.
These results demonstrate that VisU offers a fully automated, human-like approach for the mechanistic interpretation of NAMD trajectories, substantially lowering the barrier to routine photochemical analysis. By interating multimodal intelligence, VisU transcends traditional data processing and more closely emulates the working patterns of a team of computational scientists, effectively bridging the gap between raw numerical outputs and high-level chemical insight. This vision-driven framework not only redefines the protocol for analyzing complex data in excited-state dynamics simulations, but also provides a transformative blueprint for AI-augmented discovery across theoretical and computational chemistry. Ultimately, VisU represents a shift toward an agentic research paradigm by seamless combining visual reasoning and domain-specific chemical knowledge , empowering researchers to uncover hidden molecular mechanisms with unprecedented efficiency and clarity in the data-rich era.
@@ 4.2 LLM Agents and Workflow Automation
Doubao-Seed-1.6-Vision provide a substantial advantage in handling complex multimodal reasoning tasks. Leveraging its 256k-token context window, the model could accommodate multiple trajectory projections, long textual descriptions, and optional trajectory videos within a single prompt for comprehensive analysis. Context cachings ensured the automatic persistence of historical information, while structured outputs enabled efficient post-processing and downstream integration. Within VisU, these capabilities allow the Mentor to inspect multiple candidate channels concurrently, evaluate structural patterns, and generate precise, human-like mechanistic guidance.
DeepSeek-V3.2 integrates advanced autonomous tool-calling within a chain-of-thought (CoT) reasoning paradigm. In VisU, it empowers the Engineer agent to adaptively generate and execute context-aware preprocessing scripts and analysis codes, dynamically interact with the file system, and coordinate with Student modules to ensure correct task execution. These features enable the automated handling of heterogeneous simulation outputs, the flexible adaptation to previously unseen data archives, and the effecient construction of all computational codes and scripts.
@@ 5 Author contributions
Yifei Zhu: Conceptualization, Methodology, Software, Formal analysis, Investigation, Data Curation, Writing - original draft, Writing - review & editing. Jiahui Zhang: Conceptualization, Methodology, Software, Formal analysis, Investigation, Data Curation, Writing - original draft, Writing - review & editing. Binni Huang: Data Curation, Investigation, Visualization. Zhenggang Lan: Conceptualization, Supervision, Project administration, Resources, Funding acquisition, Writing - review & editing.

### v2

@@ 2.2.1 Preprocess the Input File
Figure 1: Preprocessing pipeline implemented through the cooperation of Mentor, Engineer and Student. (a) User-provided input file (test.zip) and message. (b) Generation of the preprocessing Python script using tool calling and the provided Python function templates. (c) Final preprocess.py and the resulting XYZ files. (d) Analysis of molecular structures, including the molecular structure and atomic labels of keto isocytosine shown in mol.png. (e) Six generated descriptor sets based on RICs.
@@
Clearly, as with output files generated by other NAMD packages, the raw JADE data requires extensive preprocessing before formal analysis. This preprocessing includes unzipping all output files, validating their contents, collecting the initial geometries, identifying the hopping times, and extracting the corresponding hopping geometries. Here, we emphasize that although the workflow and performance of VisU are demonstrated using keto isocytosine based on JADE outputs, the overall data-processing pipeline of VisU is, in principle, directly applicable to outputs from other NAMD codes without further modification.
@@
Rather than directly generating a complete preprocessing script in a single step, the Engineer first reasoned about the information required for subsequent analysis, including the directory structure, file naming conventions, and the locations of trajectory and geometry files. Guided by this reasoning, it selectively invoked file-system inspection tools embedded within VisU to explore the unzipped archive and extract the necessary structural information. After acquiring the workspace context, the Engineer composed a preprocessing script from a curated library of expert-validated function templates. In this sense, the overall preprocessing procedure closely resembles human expert operations. The Student then executed the script and returned error messages in cases of failure. In the present example, the program encountered a Python error (FileNotFoundError) during the first execution. The Student collected all error information and relayed it to the Engineer, which subsequently revised the script. The data-preprocessing task was completed successfully after a second iteration. Details of the error analysis and the corrected preprocessing script are provided in the SI. As discussed in the previous section, this preprocessing step produced three required output files, namely ref.xyz, init.xyz, and hopping.xyz (Figure 1.(c)).
@@
To generate molecular descriptors, VisU processed both the reference (ref.xyz) and hopping (hopping.xyz) geometries. The Student plotted a 2D molecular image (mol.png) with atomic labels in the basis of the reference geometry, as illustrated in Figure 1.(d). This image was then passed to the Mentor, powered by the Doubao-Seed-1.6-Vision model, which was carefully prompted to understand the context of the analysis workflow.
Leveraging its visual-reasoning capability, this Mentor first interpreted the image to determine whether the input molecule was compatible with the current available protocol. For the present system, VisU identified the atoms constituting the ring moiety, which provided explicit guidance for the Student to construct the corresponding RIC descriptors. Using the hopping geometries collected in hopping.xyz, the Student then built six sets of RIC descriptors, as shown in Figure 1.(e): $D_{\mathrm{ring}}$, $A_{\mathrm{ring}}$, $R_{\mathrm{ring}}$, $D_{\mathrm{eg}}$, $A_{\mathrm{eg}}$, and $R_{\mathrm{eg}}$. These descriptors formed the foundation for subsequent unsupervised channel discovery and mechanistic analysis. In parallel, the same descriptor sets were computed for all initial geometries in init.xyz.
@@ 2.2.2 Extract the Channels
Figure 2: Mentor-Student” cooperation during the first iteration of unsupervised channel discovery. Based on diagnostic plots produced by the Student, the carefully prompted visual-language Mentor (Doubao-Seed-1.6-Vision) provides targeted guidance on: (a) selection of the optimal reduced-dimensional subspace, (b) choice of clustering algorithm, (c) tuning of hyperparameters, and (d) final evaluation of the resulting clusters.
@@
Using the prepared descriptor sets, channel discovery proceeded via an iterative PCA-then-clustering analysis protocol zhuPrincipalComponentAnalysis2022a.
@@
As shown in Figure 2.(a), the first iteration of channel discovery began by applying PCA to each of the six descriptor sets to obtain low-dimensional representations of the hopping geometries. In all cases, the first two principal components contributed more than 60% of the total variance. The resulting PCA projections were then forwarded to the Mentor for assessment. In the present example, the Mentor leveraged its visual-reasoning capability to identify that the hopping points were well separated in the $D_{\mathrm{ring}}$-based reduced space, indicating the presence of several distinct channels and highlighting the important role of ring-associated dihedral motions in the nonadiabatic dynamics.
@@
The next critical step is to group similar hopping geometries into clusters. In Figure 2.(b), the Mentor inspected the distribution in the $D_{\mathrm{ring}}$ PCA plot using its visual-reasoning capability and recommended agglomerative clustering for the initial partitioning. As shown in Figure 2.(c), the Student first performed clustering with the default parameter $k=2$. Upon reviewing the resulting plot, the Mentor immediately identified that the central region was visually distinct and was not be merged with the adjacent cluster. Therefore, the Mentor instructed the Student to adjust the parameter to $k=3$, after which the clustering result aligned with expectations. Following re-clustering with the corrected parameter, the Mentor confirmed that the current clustering was accurate.
@@
A final visual inspection by Mentor, shown in Figure 2.(d), verified that the current clustering result with three clusters was reasonable, ensuring that this step aligned with expert-level visual intuition. The first discovery iteration thus concluded with three clusters: Cluster 0, Cluster 1, and Cluster 2. Among them, Cluster 2 contained only 37 structures, corresponding to less than 15% of the total dataset. This cluster was therefore treated as an independent candidate channel (Candidate Channel 0), while the remaining two clusters proceeded to a second iteration of analysis. Afterward, the same PCA-then-clustering procedure was then recursively applied to all remaining subsets. As illustrated in Figure S1, all clusters either became non-separable after three iterations or contained fewer than 15% of the structures, ultimately yielding seven nonadiabatic decay channels (Candidate Channel 0 to 6).
@@
As different data patterns appeared in these iterative steps, distinct clustering algorithms were recommended based on their distributions: including agglomerative clustering (Iteration 1), DBSCAN (Iteration 2), and K-Means (Iteration 3), as in Figure S1. Additional results are provided in the SI. These observations demonstrate that VisU can effectively handle both algorithms like agglomerative clustering, which requires only a single hyperparameter and aligns closely with visual intuition, and more complex methods such as DBSCAN, where the two parameters (eps and min_samples) are less directly interpretable from visual inspection.
@@ 2.2.3 Identify the Important Molecular Motion
At this stage, the Student performed PCA projections of both the initial and hopping geometries, submitting the resulting two-dimensional representations to the Mentor for analysis. By visually inspecting these projection plots, the Mentor determined which sets of RICs were most effective for distinguishing hopping from initial geometries. In the present example (Figure 3.(a)), the hopping and initial geometries were clearly separated in the reduced space defined by $D_{\mathrm{ring}}$, $D_{\mathrm{eg}}$, $A_{\mathrm{ring}}$, and $R_{\mathrm{ring}}$, indicating that these RIC sets captured the leading reactive coordinates in the photochemical processes.
@@ 2.2.4 Validate and Summarize Mechanisms
Figure 4: (a) Final nonadiabatic channel analysis tree. Details regarding the division and naming of different clusters can be found in the SI. (b) Averaged and (c) representative structures (hydrogen atoms omitted) for the four channels.
@@
An essential task for the Mentor is to check for chirality. For example, several candidate channels in the present system were identified as mirror counterparts based on both qualitative visual inspection and quantitative data comparison. As shown in Figure 4.(a), the Mentor merged different groups showing mirror-symmetry: Candidate Channel 1 and Candidate Channel 6 into Channel B; Candidate Channel 2 and Candidate Channel 4 into Channel C; and Candidate Channel 3 and Candidate Channel 5 into Channel D. Candidate Channel 0 exhibited no mirror-symmetric counterpart and was therefore retained as Channel A. As in Figures 4.(b) and (c), the initial seven candidate channels were finally consolidated into four distinct nonadiabatic channels: Channel A, B, C, and D.
It is important to note that different decay channels may be governed by the same nuclear motions but with varying magnitudes. For instance, Channel B, C, and D all exhibited pronounced out-of-plane motions of O9 and C10, while their direction differed. In principle, one can perform systematic pairwise comparisons of initial and hopping geometries across channels; however, this approach becomes tedious in multichannel situations. Instead, VisU relies on the Mentor to perform a holistic comparison by integrating key features, representative geometries, and dominant motions across all channels. Through direct visual inspection and cross-channel assessment, the Mentor clarifies the mechanistic similarities and differences among the nonadiabatic decay channels.
Finally, the Mentor drew the following conclusions. Channel A accounted for approximately 9% of the total hopping events, which was characterized by a dominant C9-C10 stretching motion. In contrast, the remaining three channels involved combinations of C1 puckering and O9-C10 out-of-plane motions. Among them, Channel B, Channel C, and Channel D contributed approximately 10%, 75%, and 6% of the total population, respectively. Although these three channels shared similar core motions, they differed in the directionality of the side-group motion associated with the O9-C10 fragment, as shown in Figures 4.(b) and (c). These results are consistent with our previous and recent studies zhuPrincipalComponentAnalysis2022a, liu2025automatedframeworkanalyzingstructural, and additional discussion are provided in the SI.
At the end, the complete mechanistic interpretation, including channel populations, representative structures, and visual summaries derived from the raw nonadiabatic dynamics trajectories, was compiled in a comprehensive analysis report provided in the SI. With this, the end-to-end analysis performed by VisU is complete, where was executed fully automatically without human intervention.
@@ 3 Conclusion
The powerful analysis capability of VisU is stems from the integration of two state-of-the-art large language models: Doubao-Seed-1.6-Vision, which provides expert-level visual reasoning, and DeepSeek-V3.2, which enables adaptive tool-calling and script construction. This synergy facilitates a multiagent paradigm that orchestrates a comprehensive four-module workflow: Preprocessing, Recursive Channel Discovery, Important-Motion Identification, and Validation and Mechanistic Summary.
In a representative case study, VisU successfully identified four distinct nonadiabatic channels, uncovered the corresponding leading reactive coordinates, and automatically handled mirror-symmetry merging. Furthermore, it performed cross-channel comparisons and generated a comprehensive, figure-rich academic report without human intervention. The framework demonstrated a professional-grade ability to reproduce dominant mechanistic features reported in previous work zhuPrincipalComponentAnalysis2022a, liu2025automatedframeworkanalyzingstructural. Remarkably, starting only from raw NAMD simulation outputs and minimal text prompts, the entire protocol is executed through autonomous visual inspection and expert-like guidance, powered by advanced vision-reasoning LLMs. These results demonstrate that VisU offers a fully automated, human-like approach for the mechanistic interpretation of NAMD trajectories, substantially lowering the barrier to routine photochemical analysis.
VisU demonstrates that visual reasoning is not merely a supplementary tool but a primary driver for mechanistic discovery in theoretical chemistry. This framework successfully bridges the long-standing gap between visual experience (the qualitative patterns recognized by experts) and numerical experience (the quantitative data from theoretical simulations). By integrating multimodal intelligence, VisU transcends traditional data processing and more closely emulates the working patterns of a team of computational researchers, providing professional academic analysis with the higher-level perception and reasoning required for complex scientific discovery. This vision-driven framework not only redefines the protocol for analyzing complex data in excited-state dynamics simulations, but also provides a transformative blueprint for the whole AI4Chemistry field. Ultimately, VisU represents a shift toward an agentic research paradigm by seamlessly combining visual reasoning and domain-specific chemical knowledge, empowering researchers to uncover hidden molecular mechanisms with unprecedented efficiency and clarity in the data-rich era.
@@ 4.2 LLM Agents and Workflow Automation
Doubao-Seed-1.6-Vision provides a substantial advantage in handling complex multimodal reasoning tasks. Leveraging its 256k-token context window, the model could accommodate multiple trajectory projections, long textual descriptions, and optional trajectory videos within a single prompt for comprehensive analysis. Context caching ensured the automatic persistence of historical information, while structured outputs enabled efficient post-processing and downstream integration. Within VisU, these capabilities allow the Mentor to inspect multiple candidate channels concurrently, evaluate structural patterns, and generate precise, human-like mechanistic guidance.
DeepSeek-V3.2 integrates advanced autonomous tool-calling within a chain-of-thought (CoT) reasoning paradigm. In VisU, it empowers the Engineer agent to adaptively generate and execute context-aware preprocessing scripts and analysis codes, dynamically interact with the file system, and coordinate with Student modules to ensure correct task execution. These features enable the automated handling of heterogeneous simulation outputs, the flexible adaptation to previously unseen data archives, and the efficient construction of all computational codes and scripts.
@@ 5 Author contributions
Yifei Zhu and Jiahui Zhang contributed equally to this work. Yifei Zhu: Conceptualization, Methodology, Software, Formal analysis, Investigation, Data Curation, Writing - original draft, Writing - review & editing. Jiahui Zhang: Conceptualization, Methodology, Software, Formal analysis, Investigation, Data Curation, Writing - original draft, Writing - review & editing. Binni Huang: Data Curation, Investigation, Visualization. Zhenggang Lan: Conceptualization, Supervision, Project administration, Resources, Funding acquisition, Writing - review & editing.